### v1

# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

Abstract
Human-robot teams involve humans and robots collaborating to achieve tasks under various environmental conditions. Successful teaming will require robots to adapt autonomously to a human teammate’s internal state. An important element of such adaptation is the ability to estimate the human teammates’ workload in unknown situations. Existing workload models use machine learning to model the relationships between physiological metrics and workload; however, these methods are susceptible to individual differences and are heavily influenced by other factors. These methods cannot generalize to unknown tasks, as they rely on standard machine learning approaches that assume data consists of independent and identically distributed (IID) samples. This assumption does not necessarily hold for estimating workload for new tasks. A survey of non-IID machine learning techniques is presented, where commonly used techniques are evaluated using three criteria: portability, model complexity, and adaptability. These criteria are used to argue which techniques are most applicable for estimating workload for unknown tasks in dynamic, real-time environments.
Index Terms:
human-robot teams, non-iid machine learning, workload estimation for unknown tasks
publicationid: pubid:
I Introduction
Deploying human-robot teams in uncertain environments requires robots to have a dynamic understanding of humans’ internal state. This dynamic understanding must account for real world complexities in order to enable fluid interactions between a robot and a human teammate. Workload represents how hard a person is working, and can be decomposed into workload components (e.g., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile). Incorporating accurate workload models will enhance the robot’s understanding of its human teammate, allowing the robot to manage a human’s workload level, either through intelligent task redistribution or modulation of how the robot interacts with the human. However, real world human-robot teams will inevitably encounter unseen tasks and environments.
An adaptive teaming system requires machine learning models that can estimate workload accurately in both known and unknown situations. Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days. Workload is also influenced by factors, such as experience, stress, and fatigue [1]. Generalizing to novel situations while also accounting for these factors is difficult for traditional machine learning models, as these standard approaches assume data consists of independent and identically distributed (IID) samples [2]. However, this assumption does not hold for estimating workload for new tasks.
Consider first response robotics. Urban firefighters will work alongside robots to perform a wide range of tasks [3, 4]. Firefighters have historically used robots to gather aerial imagery of buildings, but more advanced control algorithms will enable the use of similar robots to extinguish fires in difficult to access locations [4]. This new tasks require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
Prior work established that standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., [1, 5, 6]), but there is no widely accepted approach that can generalize to unknown tasks [7]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of workload components; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1].
This dynamic relationship between physiological signals and an individual’s underlying workload is exemplar of the variability of human behavior, which makes the development of machine learning models that generalize in the real world difficult. Gathering the necessary data to develop these models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of human activity within real world domains. Similar complexity is seen across human-robot interaction (HRI) application domains [8, 9, 10]. The real world variability of human behavior and activity violates the IID assumption, requiring non-standard machine learning techniques (e.g., continual learning, meta-learning) to successfully overcome these challenges [11].
The manuscript surveys applicable non-IID machine learning methods. Specifically, the non-IID methods are evaluated to assess their viability for developing models of the human’s internal state (e.g., workload estimation) that can accommodate real world variability. Workload estimation for unknown tasks is an exemplar of the challenge in constructing machine learning models for HRI application domains that are characterized by non-IID data. Section II discusses the fundamental of objective workload estimation and distribution shifts that characterize non-IID machine learning. Section III lays out the notation and evaluation criteria. Section IV evaluates a wide range of non-IID machine learning methods, and Section V discusses the most applicable methods for workload estimation for unknown tasks. Finally, Section VI concludes and presents potential future work.
II Background
Workload is a complex, dynamic, individual-specific, non-linear construct, and developing models that can account for these complexities is a non-trivial task. Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown [7] [12]. Modern techniques (e.g., meta-learning) approach machine learning from a different perspective and possess the potential to build more flexible workload estimation models.
Humans have limited resources for processing information, making decisions, and dealing with physical stress. Formally, workload is the degree of activation of a finite pool of resources while performing a task [7] that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience). Workload has been shown to vary across individuals, across tasks, and over time [13] [14]. Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical [15], and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile [16].
Workload estimation methods either rely on subjective (e.g., NASA Task Load Index [17], In-Situ Surveys [18]) or objective metrics. Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports [19] [20]. Robots operating in dynamic environments need high-frequency workload measures to adequately adjust to their human teammates’ current state. Objective estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., [21, 22, 23, 24, 25]).
Human-robot teams that are deployed in unstructured, dynamic environments cannot rely on environmentally embedded sensors (e.g., built-in cameras), or complex sensor systems (e.g., electroencephalogram (EEG)). Thus, wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently. Each workload component corresponds to a different set of physiological metrics, though some metrics may correspond to multiple components. A good example is pupil diameter, which is positively correlated with cognitive and visual workload [16]. Robots need a reliable estimate of the distribution of workload across the components, as tasks will change over time, and differing tasks impact different components.
Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability [16]. Sensitivity represents an algorithm’s ability to reliably detect workload levels (i.e., $\geq$ 80% accuracy or $\leq$ 5 root mean squared error (RMSE)). Diagnosticity refers to the capability of detecting different types of workload. Suitability refers to an algorithm’s ability to assess the complete overall workload state. The last metric, generalizability, refers to the algorithm’s ability to generalize across individuals and tasks. This metric defines “generalizing across tasks” as an algorithms capacity to asses the complete overall workload state (i.e., each workload component) and defines “generalizing across individuals” as an algorithms capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., [26, 27, 28, 29]). Many methods are either suitable (e.g., [30, 31, 32, 33]) or sensitive (e.g., [34, 35, 5, 36]), but very few algorithms are diagnostic. Further, little work has been done to evaluate the generalizability of existing workload estimation methods.
Prior work tends to focus on either generalizing across individuals or generalizing across tasks. Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., [37, 24, 38, 39]); however, generalizing across tasks, either known or unknown, is rarely evaluated. Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms [40]. Much of this work evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., [40, 41, 42, 43, 44]). Thus, these algorithms are sensitive, but not generalizable. Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., [45, 46, 47]); thus, failing to meet the generalizability criterion.
A support vector machine-based workload estimation algorithm that uses Recursive Feature Elimination to optimize the input features was shown to estimate cognitive workload on arbitrary tasks [48]. This method achieved $<$ 1.0 mean squared error on cross-task workload estimation. Similar methods were developed by Guan et al. [23] and Zhou et al. [49]. Additionally, deep learning techniques (e.g., attention mechanisms) that estimate workload have shown promise [50]. However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from a static sensor (i.e., EEG [51, 52, 53]). Thus, none of the methods are viable for use in dynamic, uncertain environments.
Real world human-robot teams will inevitably encounter unique scenarios when deployed. An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences [14]. Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples [2]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1]. This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques [2].
III Notation and Evaluation Criteria
Dynamic human-robot teams will rely on machine learning methods to enable a robot to quantify different aspects of their human teammate (e.g., workload). A major challenge is that these teams will continuously encounter new scenarios that the robot’s underlying machine learning methods cannot account for. These new scenarios constitute a meaningful distribution shift from conventional machine learning method’s training dataset (i.e., violate the IID assumption). Specifically, a distribution shift indicates scenarios where training and testing data come from different distributions. Understanding the difference between non-IID machine learning methods is key for building human-robot teams that can handle the variability of real world problem domains.
The workload estimation and the machine learning literature both cover a large spectrum of work. Naturally, there are terms that frequently occur in both fields which require disambiguation. The term task is used in the workload literature to refer to a job an individual may perform (e.g., lifting boxes, mental calculations), but the machine learning literature uses the term to refer to a learning task for a machine learning model. The term domain is often used to characterize the input space for a machine learning problem. For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather [54]. Domain may also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
The term learning task will refer to instances of a machine learning problem. The term task, without any adjectives, will refer to a task conducted by a human. The term domain will refer to different input spaces that a machine learning algorithm may operate on. The term application domain will refer to applications where human-robot teams may be deployed.
Definitions of domain and learning task, in the context of distribution shift, have long been established [55]:
Definition III.1
(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},…,x_{n}\}\in\mathcal{X}$.
Definition III.2
(Learning Task) A learning task $L$ is a two-tuple $(\mathcal{Y},f(x))$ associated with a specific domain $D$. $\mathcal{Y}$ is the label space of $D$ and $f(x)$ is an objective predictive function for $D$, which is learned from the training data. $f(x)$ can be written as a conditional probability distribution $P(y|x)$.
Grounding this notation in workload estimation helps to provide an intuitive explanation of the different variables. A particular workload estimation instance represents the domain, $D$, which is characterized by the set of physiological metrics, $\mathcal{X}$. The marginal distribution, $P(X)$, represents the probability distribution over the physiological metrics’ possible values.
A learning task, $L$, is a machine learning problem that is characterized by a label space $\mathcal{Y}$ and a predictive function $f(x)$, corresponding to a single domain $D$. The label space represents a set of possible workload values, and the predictive function produces a workload value, $y_{i}\in\mathcal{Y}$ given a set of physiological metrics, $x_{i}\in\mathcal{X}$. A predictive function can take the form of a conditional probability distribution $P(\mathcal{Y}|\mathcal{X})$, which represents the probability of assigning a workload value to a given set of physiological metrics.
It is important to note that changes to either $D$, $\mathcal{Y}$, or $f(x)$ constitute a new learning tasks. Changes to a domain (i.e., domain shift) are intuitive, as they represent a distribution shift of physiological metrics brought on by performing a task under different working conditions (e.g., weather). Changes to a label space (i.e., label shift) are characterized by a different range of possible workload values or distribution across workload components, typically brought on by performing a new task. Changes to the predictive function $f(x)$ are less intuitive and more easily explained via task recognition.
A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., (1, 2, and 3)), where each integer represents a unique task. A similar task recognition problem can be defined using the same metrics and the same label space, but where each integer represents the task difficulty (i.e., easy, medium, or hard). These two task recognition problems are semantically different, but the underlying machine learning method does not inherently recognizes these differences; therefore, uniqueness of these two learning tasks cannot be fully specified by the domain and label space. The changes in how the resulting predictive functions, $f(x)$, map the physiological metrics to the output values is a key characteristic in differentiating these two learning tasks.
Numerous methods have been applied to the non-IID machine learning problems, each characterized by the number of distributions and distribution shift type (see Fig. 1). Three evaluation criteria are applied to the presented methods: (1) Portability (2) Model Complexity (3) Adaptability.
Portability refers to the number of distributions used to train and test a model [56]. Learning tasks may consist of multiple source domains (e.g., rural hospital data vs. urban hospital data). A straightforward means of combining information from multiple domains is to construct a single joint distribution; however, data in some domains may be of higher quality than others. Weighting high-quality data over low-quality data is appropriate to maximize performance [56]. Additionally, the separation of source domains may be artificial. Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks [57]. Portability can have four values: one-to-one, one-to-many, many-to-one, and many-to-many. One-to-one refers to an algorithm that transfers knowledge between one source and one target domain. One-to-many transfers knowledge from one source domain to many target domains, and many-to-one transfers knowledge from multiple source domains to a single target domain. Many-to-many knowledge transfer occurs between multiple sources to multiple target domains simultaneously.
Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of [2].
Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful [58]. Machine learning models with more parameters require higher volumes of training data. Early statisticians built simple linear models using hundreds of examples [58], whereas modern deep learning architectures may require tens of billions of examples [59]. Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as more complex models are often required to solve more complex problems. Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for machine learning model development; thus, smaller models are preferred. An example of low model complexity is a random forest model that contains a few hundred decision trees, where the depth of each tree is dataset-dependent; thus, these models typically contain on the order of thousands of parameters [60]. An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters [22], but may contain on the order of millions of parameters. An example of high model complexity is deep neural networks (e.g., convolutional networks, long short-term memory networks (LSTM)), also referred to as deep learning architectures. Deep learning architectures can contain anywhere from tens of millions [61] to over 100 billion parameters [59].
The last metric, Adaptability, refers to the amount of data required to update a machine learning model to solve new problems [62] [63]. Many non-IID machine learning methods seek to account for distribution shifts by updating existing machine learning models’ parameters so that the model can solve new learning tasks. Adaptability is the number of test data points necessary for the updated model to make accurate predictions for new learning tasks, where the goal is to minimize the required data. A machine learning method has no adaptability if it requires the same amount of training data in the source and target domains. Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance [58]. Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain [63]. A 50% reduction in data required for updating a model is non-trivial, and may still constitute a large dataset. An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts [62]. It is important to note adaptability is application domain-specific, as the learning tasks are fundamentally different.
IV Non-IID Machine Learning Methods
There are several non-IID machine learning methods that can account for different distribution shift types, including: transfer learning, domain adaptation, test-time training, continual learning, domain generalization, and few-shot learning. Each method corresponds to a specific distribution shift type and the number of distributions considered (see Figure 1). There are two primary distribution shifts: (1) Domain shift, and (2) Label shift. A domain shift indicates a distribution shift of the input features, and label shift indicates a shift of the output values. Label shifts can be characterized by either a new distribution of output values, or a new set of values.
Standard IID methods only account for data from a single distribution, which can lead to performance degradation in the presence of distribution shifts [2]. Non-IID methods seek to account for these shifts by intelligently incorporating information from multiple distributions (i.e., domains). Further, these methods either account for exactly two distributions or from more than two distributions, where the boundaries between distributions may be real or artificial. The variability of human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments. However, the data can be decomposed or re-organized such that any of the presented techniques can be applied. Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks [64], or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models [63].
Many of these methods have been successfully applied to task recognition using wearable sensors (e.g., [65, 66, 67, 68]), which shares core commonalities with workload estimation. Thus, these methods provide insight into the capabilities of the underlying machine learning models to account for distribution shifts in relevant application domains. Prior work has primarily focused on evaluating standard IID methods’ ability to generalize to new situations, where many methods fail to generalize to new tasks (e.g., [40, 43, 48, 47]). Relying on standard IID methods ignores the distribution shifts of the physiological signals and workload components as tasks change. No prior work modeled workload estimation for unknown tasks as a non-IID problem. An appropriate non-IID machine learning method for the workload estimation for unknown tasks must be identified, so that the resulting workload estimation algorithm can account for distribution shifts across tasks.
IV-A Domain Adaptation
Domain adaptation considers the setting where training data and test data are characterized by the same features, but are sampled from different distributions.
Definition IV.1
(Domain Adaptation [69]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$.
This distribution shift type is called a domain shift, or a covariate shift. For example, consider multiple datasets with images of the same objects [70]. Changes in the images’ lighting, background, or orientation represent meaningful difference in the input features and are prototypical examples of a domain shift. Domain adaptation is a non-IID machine learning method that seeks to minimize performance degradation in the presence of domain shift. It is important to note that domain adaptation techniques focus on distribution shifts between one source and one target domain (i.e., two distributions), as shown in Figure 1.
Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training in order to perform well [55]. Further, there are two forms of domain adaptation: supervised [70] and unsupervised [61]. Supervised domain requires that labels be available for $L_{T}$, and there are no labels available for $L_{T}$ for unsupervised domain adaptation. There tends to be significantly more data available in source domains than the target domains, and using labeled data from both domains typically results in overfitting to the source distribution [70]. Thus, the majority of prior work has focused on unsupervised domain adaptation.
Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., [71, 72, 65, 73, 74]), workload estimation (e.g., [49, 75, 76]), and emotion recognition (e.g., [77, 78, 79, 80]). The vast majority of these applications use adversarial learning techniques [66, 75, 80], though self-supervised and feature representation learning strategies have also been moderately successful [77].
Adversarial domain adaptation consists of three model types: (1) feature extractor, (2) domain discriminator, and (3) label predictor. Seminal work integrated adversarial learning and domain adaptation as a mini-max game [81]. This work demonstrated that domain adaptation can be achieved by jointly optimizing a feature extraction model to maximize the difference of latent features from source and target domain data, via a domain classifier, and to minimize the error in predicted values for learning tasks, via a label predictor. This joint optimization enables the effective learning of a useful latent feature representation. Prior work demonstrated that replacing the discriminative feature extraction models with generative feature extraction models is also an effective means of domain adaptation (e.g., [82, 83, 84]).
Most domain adaptation techniques exhibit one-to-one portability [84], but multi-target domain adaptation (i.e., one-to-many portability) is an active research area [85]. Domain adaptation techniques primarily rely on deep neural networks [49, 75, 77], consisting of millions of parameters; thus, exhibiting a high model complexity. The adaptability of domain adaptation varies. Some approaches require one example in the target domain [86], while others require thousands [66]. Further, access to source and target data during training is a hard requirement for many domain adaptation techniques, which makes them impractical for real world HRI. One-to-one portability, high model complexity, and low adaptability are undesirable properties; thus, domain adaptation techniques are not viable for workload estimation of unknown tasks.
IV-B Transfer Learning
Transfer learning aims to improve the performance of a machine learning model for a target domain by transferring knowledge from a different, but related source domain.
Definition IV.2
(Transfer Learning [63]) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$.
Transfer learning is characterized a label shift [87], which can refer to a shift in the relative distribution of output values (e.g., disease prevalence in a rural hospital vs. urban hospital), or a complete change in the labels available. Transfer learning techniques are capable of transferring knowledge from a model trained to identify everyday objects (e.g., cats, planes) in images to a model performing medical image analysis (e.g., ultrasounds, x-rays) [88]. Similar to domain adaptation, transfer learning has been widely applied to HRI (e.g., human task recognition [89, 90], emotion recognition [91, 92, 93, 94]). There are many transfer learning techniques, which can be categorized based on how knowledge is transferred: instance-based (e.g., [95, 96, 97]), feature-based (e.g., [98, 99, 100]), and parameter-based (e.g., [101, 87, 68, 102]). Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques [103].
Parameter-based methods have two phases: pre-training and fine-tuning. Pre-training trains a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$. Fine-tuning trains the same model on data from the target domain, $D_{T}$, but only updates a subset of the parameters. This two-step training procedure enables the network to learn a useful latent representation during pre-training, and a task-specific transformation of that representation during fine-tuning. This approach has been successful in computer vision [87], robotics [104], and task recognition (e.g., [105, 106, 107, 108]).
Many of these methods use deep learning architectures to transfer knowledge across source/target domain pairs, requiring thousands of examples to be successful. Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability [63] [59]. Zhuang et al. evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability [63]. Similar findings exist in transfer learning for human activity recognition [105]. Similar techniques were also developed to solve a wide range of natural language problems based on only a few examples [59]. This high adaptability is in direct contrast to the previously discussed transfer learning techniques, where the primary difference is the model complexity. A transformer deep learning architecture was trained to predict the next word in a sentence [59]. This network was the foundation of several downstream networks that performed tasks, including machine translation, question answering, and arithmetic word problem-solving. The architecture consisted of 175 billion parameters, requiring significantly more data to train parameter-based techniques that rely on smaller networks (i.e., tens of billions of data points). These results suggest that a higher model complexity can directly improve portability and adaptability.
A core limitation of workload estimation is data volume, as gathering large volumes of wearable sensor data is logistically impractical; thus, model complexity is constrained. It is also desirable that workload estimation algorithms exhibit many-to-many portability, as models may need information across multiple individuals and tasks. Thus, transfer learning is not a viable option for workload estimation of unknown tasks.
IV-C Test-time Adaptation
Domain adaptation techniques learn to anticipate domain shifts by simultaneously training on $D_{S}$ and $D_{T}$; however, target domain data may not be available during training. Test-time adaptation techniques seek to overcome this limitation.
Definition IV.3
(Test-time Adaptation [109]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$.
Test-time adaptation techniques seek to overcome domain shifts by separating the training process into two distinct phases: (1) training on $D_{S}$ data, and (2) updating that model with $D_{T}$ data. This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 [109].
Prior work applied test-time adaptation techniques to egocentric task recognition [110] and adaptive policy optimization for assistive robots [111]. However, test-time adaptation techniques have not been widely applied to HRI application domains. Nevertheless, these techniques hold promise because they overcome the limitation of training on source and target domain data simultaneously. These techniques can be categorized into as either self-supervised, or entropy-based.
Self-supervised techniques use a dual-output neural network and a two-phase training procedure [109]. Labeled data from the $D_{S}$ undergoes some deterministic transformation (e.g., image rotation). The neural network is trained on this transformed data to predict both the original label (e.g., image class) and the corresponding transformation (e.g., rotation angle). The unlabelled data from $D_{T}$ undergoes the same deterministic transformation. The shared feature extraction model’s parameters are updated based on the output of the self-supervised proxy learning tasks. This two-phase procedure decouples training from adaptation and enables the model to adapt at test-time. Many techniques take inspiration from this work (e.g., [112, 113, 114]); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., [115, 116, 117]). Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy [118] is calculated for each batch. The adaptation phase updates the shared feature extractor to minimize the average entropy over all batches [119].
Fundamentally, test-time adaptation is domain adaptation without access to target domain data during training. Therefore, test-time adaptation exhibits one-to-one portability, high model complexity, and low adaptability. It is difficult to definitively state whether these techniques are applicable to workload estimation for unknown tasks, as the relationship between model complexity and adaptability for these techniques is unknown. Prior work has primarily focused on deep learning architectures, but these techniques are general enough to apply to shallow neural networks. It is possible that smaller networks may possess sufficient adaptability to account for real world variability, but the efficacy of test-time adaptation on these networks has not been thoroughly evaluated.
IV-D Continual Learning
Continual Learning techniques train a model to perform a sequence of non-IID learning tasks, where each task is characterized by a unique domain shift (see Figure 1).
Definition IV.4
(Continual learning [120]) Given a sequence of domains $D$ = $D_{s_{1}},...,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},…,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$.
These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks [121]. Continual learning is also referred to as lifelong learning, sequential learning, or online learning [122]. There is a wide range of techniques, including knowledge distillation (e.g., [123, 124, 125]), memory-based methods (e.g., [126, 127]), and parameter isolation (e.g., [128, 129]).
Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., [130, 131, 132, 133]), affective computing (e.g., [8, 9, 134]), and adaptive human-robot interactions (e.g., [8, 135, 136]).
Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary [137]. Thus, continual learning techniques can be trained offline as the number of tasks is known a priori. These techniques are referred to as multi-head techniques, as separate models are trained for each learning task [138]. However, knowing all the tasks ahead of time is not always realistic. Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online [138]. These techniques are referred to as single-head techniques, since a single model is trained for all tasks.
Learning tasks within a sequence can either be independent or dependent. Independent learning tasks are separate and may not share similarities. Dependent learning tasks represent a sub-sequence of learning tasks that must be executed in a particular order, where learning task $L_{i}$ holds key information about learning task $L_{i+1}$. Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques [139, 140].
The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes [136], shallow neural networks [127], and deep networks [122]). Many techniques exhibit low adaptability, requiring thousands of examples to perform well [122]. Portability is challenging to define for continual learning, as it exhibits many-to-many portability, but does so only in a sequential fashion. Many-to-many portability is desirable, but the sequential nature is an unnecessary constraint. Additionally, some continual learning methods only optimize performance on the latest learning tasks, while other optimize performance on all prior learning tasks. Portability and adaptability are key qualities required to estimating workload successfully for unknown tasks; thus, continual learning is a subpar option.
Figure 2: Comparison of applying the technique Model Agnostic Meta-learning [141] to: (a) Domain Generalization, sub-problems experience a domain shift, and (b) Few-shot Learning, sub-problems experience a domain and label shift.
IV-E Domain Generalization
Domain generalization considers the setting where training data from multiple source domains is characterized by the same features that are sampled from different distributions [142]. Specifically, domain generalization is the process of transferring knowledge acquired from several related source domains and applying it to previously unknown domains [143], as indicated in Figure 1.
Definition IV.5
(Domain Generalization [142]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$.
The goal is to leverage unique cross-domain information to improve the machine learning model’s generalization. These methods assume the construction of a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains. A key feature of domain generalization is the lack of labeled target domain samples. Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., [144, 145, 146]), data manipulation (e.g., [147, 148, 149]), and learning strategy (e.g., [150, 151, 152]).
The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned [153]. Domain alignment techniques measure this difference in many ways, including minimizing moments [154], contrastive loss [155], KL divergence [156], or maximum mean discrepancy [157]. Prior work combined domain alignment metrics with domain-adversarial learning [157]. The primary difference between domain-adversarial techniques for domain generalization and domain adaptation is that the source/target discriminator must differentiate between multiple source domains.
Data manipulation techniques seek to artificially enhance the size and diversity of a dataset via data augmentation or generation. Data augmentation performs deterministic transformations over existing data to create new examples [158]. However, data augmentation is not domain-agnostic, as useful transformations are specific to an application domain. Data generation uses generative machine learning models (e.g., Variational Autoencoders [147], Generative Adversarial Networks [148]) to create completely new training examples. These techniques are domain-agnostic, but the underlying models are hard to train and require a high training data volume [159]. These techniques cannot be applied to workload estimation, as generating realistic physiological metrics that correctly correspond to the workload values is non-trivial.
Learning strategy-based domain generalization techniques seek to improve cross domain generalization through novel training methodologies, including: 1) Self-supervsied learning 2) Ensemble learning, and 3) Meta-learning. Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation (see Section III.C). Ensemble learning is the process of training multiple machine learning models with varying weight initializations or training data splits, then aggregating the models’ outputs to make predictions. These extensively studied techniques are notoriously computationally expensive [160] and are not viable solutions for real-time HRI domains.
Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains [141]. Broadly, meta-learning is ‘learning to learn’ [161]. Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) [141]) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
A meta-model, parameterized by $\theta$, is trained to predict an initialization for the parameters of another model, $\phi_{i}$, that performs the learning task $L_{i}$. Both $\theta$ and $\phi_{i}$ typically represent a neural network’s weights, but can also represent a subset of the network’s weights, or the another model’s parameters. Each source domain $D_{i}$ and learning task $L_{i}$ is deconstructed into a collection of mini-datasets. These mini-datasets are defined by the support set, $S_{i}$, and the query set, $Q_{i}$, which are small training and test sets, respectively. The support set consists of a few (e.g., 1, 5, 10) examples that the meta-model uses to predict the labels for the query set. The task-specific model’s error is twice differentiated and backpropagated through both models to update $\theta$. This problem formulation constructs the learning process, such that the model learns how to adapt its parameters based on a few examples. This technique is popular in meta-learning [161], and is the foundation for many techniques (e.g, [162, 163, 164, 165, 166]). An alternative approach learns the parameters of an optimizer. An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples [167]. The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer [168].
The majority of domain generalization techniques exhibit many-to-many portability by construction. Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures [169]. Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation [170], and others only require a few examples [171]. MAML and other meta-learning techniques exhibit high adaptability and are general enough to apply to shallow neural networks; thus, exhibiting a low model complexity. Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times [162].
IV-F Few-shot Learning
Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain [62].
Definition IV.6
(Few-shot Learning [62]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$.
Few-shot learning is often characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1. Prior work primarily focused on applying meta-learning techniques to solve few-shot learning problems. These techniques overlap significantly with meta-learning for domain generalization. Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML [141], Prototypical network [172], Conditional Neural Processes [173])). Meta-learning techniques broadly fall into three categories: 1) parameter-based 2) metric-based 3) probabilitstic. Parameter-based meta-learning algorithms were discussed in Section III.G. An illustrative example demonstrating the primary difference between meta-learning for domain generalization and few-shot learning is provided in Figure 2.
Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared [64, 174, 175]. Two neural networks are trained to encode data, from the source and target domain, into a latent feature space. One neural network, $g_{\theta}$, encodes the support set (i.e., source domain data), and a second, $f_{\theta}$, encodes the query set (i.e., target domain data). Typically, multiple support set data points are passed through the networks simultaneously to enable the calculation of a range of complex distance metrics.
(a) Conditional Neural Process
(a) Conditional Neural Process
(b) Latent Neural Process
Figure 3: A comparison of two basic neural process types: (a) conditional neural processes and (b) latent neural process. Green boxes represent input and output values, blue boxes represent a neural network, yellow boxes represent aggregated latent features, and the pink circle represent an aggregation function. This figure is adapted from Figure 3 in [176].
Matching networks perform a pairwise comparison between latent features using cosine similarity [64]. Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space [172]. Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks [172]. Relation networks decompose the problem into separate modules: an embedding module and a relation module [177]. The embedding module is a neural network designed for feature extraction and the relation module is a non-linear classifier trained on the latent feature representation. Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class [178]. This technique maximizes the distance between subspaces by using a Grassmannian geometry-based objective function. Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g, [179, 180, 181]).
Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process [182])) and a neural network. Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets [176]. Probabilistic meta-learning techniques fall into two broad categories: Neural Processes [183], and Deep Kernel Learning [184].
Neural processes pass the data through a neural network, parameterized by $\theta$, to learn a latent representation. Each support set data point is passed through an aggregation function, $a$, to calculate the latent feature, $z$. This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes [176]). Another neural network, parameterized by $\phi$, uses $z$ and one or more data points (e.g., one-shot vs. few-shot) from the target domain to learn the parameters of a Gaussian Process. There are two main neural process types: latent neural processes, and conditional neural processes, as shown in Figure 3.
Latent neural processes construct a stochastic variable using latent features, which typically takes the form of a Gaussian distribution parameterized by $\mu$ and $\sigma$ [183]. However, by modeling these correlations the underlying predictive model becomes intractable and must be approximated (e.g., Monte Carlo Sampling [185], Variational Inference [183]). Conditional neural processes do not construct a stochastic variable using latent features [173], instead they make the assumption the stochastic process can be fully described by these encoded values (i.e., factorized Gaussian distribution [176]). This assumption makes the underlying predictive model analytically tractable, but ignores any correlations that may exist between latent features. Several neural network architectural components have been incorporated into conditional neural processes with varying degrees of success (e.g., [186, 187, 188, 189, 190, 191, 192, 193]).
Deep kernel learning techniques takes an alternative by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, effectively training the network to learn a custom kernel function for that Gaussian process [184]. Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks [194]. This technique learns over a large number of small, but related, examples that train the underlying machine learning model to approximate a prior distribution to transfer knowledge between learning tasks. Deep kernel transfer quantifies uncertainty by maintaining a distribution over parameters of the neural network, which can become computationally expensive for larger networks. Pólya-Gamma augmentation and softmax approximation were explored to overcome this limitation by maintaining a distribution over functions [195]. Deep mean function learning has also been applied as a meta-learning technique, where a neural network is used to learn a custom mean function for a Gaussian process [196]. These and similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., [156, 197, 198, 199, 200]). However, recent results demonstrate that the original deep kernel learning algorithm tends to produce unreliable uncertainty estimates in some practical application domains [201]. Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (e.g., [201][202])
Few-shot learning techniques exhibit many-to-many portability and high adaptability. Most few-shot learning techniques were developed using computer vision benchmarks (e.g., [203, 172, 64]), which requires deep learning architectures. However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model [11]. Therefore, it stands to reason that model complexity is directly related to the learning task’s complexity. Meta-learning techniques have been proven to be successful solutions to both few-shot learning and domain generalization. Further, these techniques exhibit desirable properties for all three presented criteria. Therefore, meta-learning presents a promising option for develop workload estimation for unknown tasks solutions and merit future investigation.
V Discussion
Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are of particular importance to each task. Understanding the distribution across workload components is critical to adapting interactions and task allocations to maximize the performance of a human-robot team. Quantifying uncertainty helps enumerate this distribution by contextualizing component-specific workload estimates, as components may rely on overlapping metrics (e.g., fine-motor and tactile).
Domain adaptation and transfer learning are popular methods, but require high model complexity to achieve desired performance levels. These methods are also best suited to one-to-one portability problems with low adaptability. Specifically, these techniques are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains and where all data is gathered prior to training. These techniques do not adapt machine learning models in the moment; thus, are not viable for performing workload estimation for unknown tasks in real world application domains. Test-time adaptation overcomes this particular constraint, but still exhibits one-to-one portability and low adaptability. Existing test-time adaptation techniques applied to smaller neural networks may exhibit higher adaptability, but further investigation is required.
Continual learning has been frequently proposed as a potential solution to non-IID machine learning problems within HRI, but are not applicable to workload estimation for unknown tasks due to their poor portability. Incorporating information for an indeterminant number of tasks, while avoiding performance degradation on all prior tasks are additional constraints that likely make the machine learning problem unnecessarily difficult. These techniques learn how to transform the model between task pairs, whereas other techniques (e.g., domain generalization, few-shot learning) learn to transform from a common prior. Consider an application domain with $N$ tasks. Continual learning must learn $N!$ task-pair independent updates, as the tasks’ order influences the learning process. Domain generalization and few-shot learning techniques only learn $N$ updates, as they can trace each task to a single common prior. These additional portability-related constraints make continual learning a subpar option.
Domain generalization and few-shot learning exhibit desirable properties for all three criteria: high adaptability, one-to-one portability, and low model complexity. These techniques are explicitly designed to update a machine learning model based upon limited and highly variable data, and are flexible enough change the model complexity based upon the learning task’s complexity. Multiple domain shifts from a common prior is the most directly related metaphor for the variability seen in human behavior; thus, domain generalization is the most suited for scenarios where there is no label shift (e.g., workload estimation). Few-shot learning may be more applicable to scenarios where labels change between learning tasks or application domains (i.e., task recognition, affect recognition). The presented meta-learning techniques are general enough to apply to either machine learning problem.
Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem [1, 204, 21]. Only a limited number of techniques (e.g., probablistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks. Probablistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
A key research question is if existing workload estimation datasets capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques can accurately estimate workload for unknown tasks. It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial. These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID and non-IID tasks that these techniques can handle. Understanding how probablistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real world viability. Additionally, analyzing how the differences between tasks, human-robot interactions, robot capabilities further informs how these techniques can be successfully deployed. Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown tasks; thus, merit further investigation.
VI Conclusion
Developing models that can estimate the human’s internal state is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments. The variability of these environments is characterized by varying human-robot teaming dynamics, diverse tasks executions, and different operating conditions for the team itself. A robot seeking to adapt its behavior to best assist its human teammate must accurate the human’s internal state accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible. Non-IID machine learning, particular domain generalization and few-shot learning, hold promise in developing models that can account for these varying dynamics.
Acknowledgments
The graduate students have been supported by National Aeronautics and Space Administration through University Space Research Association aware 904186092. The contents are those of the authors and do not represent the official views of, nor an endorsement, by the National Aeronautics and Space Administration.
The presented work was partially supported by ONR grants N00024-20-F-8705 and N00014-21-1-2052. The views, opinions, and findings expressed are those of the authors and are not to be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.

### v2

@@
-Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift
+A Survey of Machine Learning for Estimating Workload: Considering Unknown Tasks
 Abstract
-Human-robot teams involve humans and robots collaborating to achieve tasks under various environmental conditions.
-Successful teaming will require robots to adapt autonomously to a human teammate’s internal state.
-An important element of such adaptation is the ability to estimate the human teammates’ workload in unknown situations.
-Existing workload models use machine learning to model the relationships between physiological metrics and workload; however, these methods are susceptible to individual differences and are heavily influenced by other factors.
-These methods cannot generalize to unknown tasks, as they rely on standard machine learning approaches that assume data consists of independent and identically distributed (IID) samples.
-This assumption does not necessarily hold for estimating workload for new tasks.
-A survey of non-IID machine learning techniques is presented, where commonly used techniques are evaluated using three criteria: portability, model complexity, and adaptability.
-These criteria are used to argue which techniques are most applicable for estimating workload for unknown tasks in dynamic, real-time environments.
-Index Terms:
-human-robot teams, non-iid machine learning, workload estimation for unknown tasks
-publicationid: pubid:
-I Introduction
+Successful human-robot teaming will require robots to adapt autonomously to a human teammate’s internal state, where a critical element of such adaptation is the ability to estimate the human’s workload in unknown situations.
+Existing workload models use machine learning to model the relationship between physiological signals and workload.
+These methods often struggle to generalize to unknown tasks, as the relative importance of various physiological signals change significantly between tasks.
+Many of these changes constitute a meaningful shift in the data’s distribution, which violates a core assumption made by the underlying machine learning approach.
+A survey of machine learning techniques designed to overcome these challenges is presented, where common techniques are evaluated using three criteria: portability, model complexity, and adaptability.
+These criteria are used to analyze each technique’s applicability to estimating workload during unknown tasks in dynamic environments and guide future empirical experimentation.
+keywords:
+human-robot teams, workload estimation for unknown tasks, machine learning under distribution shift
+1 Introduction
 Deploying human-robot teams in uncertain environments requires robots to have a dynamic understanding of humans’ internal state.
-This dynamic understanding must account for real world complexities in order to enable fluid interactions between a robot and a human teammate.
-Workload represents how hard a person is working, and can be decomposed into workload components (e.g., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile).
-Incorporating accurate workload models will enhance the robot’s understanding of its human teammate, allowing the robot to manage a human’s workload level, either through intelligent task redistribution or modulation of how the robot interacts with the human.
-However, real world human-robot teams will inevitably encounter unseen tasks and environments.
+This dynamic understanding must account for real-world complexities in order to enable fluid interactions between a robot and a human teammate.
+Workload represents how hard a person is working, and can be decomposed into workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile).
+Incorporating machine learning-based workload models that accurately capture the contribution of each component will enhance the robot’s understanding of its human teammate, allowing it to respond appropriately to undesirable workload levels (i.e., overload (OL) or underload (UL)) through intelligent modulation of how the robot interacts with the human.
+These adaptive interactions enable more collaborative human-robot teaming dynamics and facilitate long-term collaboration.
+However, real-world human-robot teams will inevitably encounter novel (i.e., unknown) tasks.
+These tasks present unique challenges for the machine learning models that underpin modern workload estimation algorithms that rely on physiological signals, as unknown tasks frequently exhibit unique data characteristics that the model was not trained to account for.
 An adaptive teaming system requires machine learning models that can estimate workload accurately in both known and unknown situations.
-Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days.
-Workload is also influenced by factors, such as experience, stress, and fatigue [1].
-Generalizing to novel situations while also accounting for these factors is difficult for traditional machine learning models, as these standard approaches assume data consists of independent and identically distributed (IID) samples [2].
-However, this assumption does not hold for estimating workload for new tasks.
+Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days (Longo et al., 2022; Wickens et al., 2004).
+Workload is also influenced by factors, such as experience, stress, and fatigue (Heard et al., 2019).
+Developing machine learning-based workload estimation models that generalized to novel situations while also accounting for these factors is difficult, as these traditional approaches assume data consists of independent and identically distributed (IID) samples (Murphy, 2023, chap.
+19).
+The IID assumption states that a machine learning model’s training and testing data (e.g., real-world data) are drawn from the same probability distribution.
+This assumption implies that the training data is sufficiently representative of real-world data, which does not necessarily hold when estimating workload for novel, unknown tasks.
+Prior work established that workload models built using standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., Heard et al., 2019; Kaczorowska et al., 2021; Manawadu et al., 2018), but there is no widely accepted approach that can generalize to novel, unknown tasks (Longo et al., 2022).
+Generalizing across tasks is difficult due to a shift in the relative balance of workload components, where certain physiological signals and workload components are of particular importance to a given task and the magnitude of that significance varies between tasks (Heard et al., 2019).
 Consider first response robotics.
-Urban firefighters will work alongside robots to perform a wide range of tasks [3, 4].
-Firefighters have historically used robots to gather aerial imagery of buildings, but more advanced control algorithms will enable the use of similar robots to extinguish fires in difficult to access locations [4].
-This new tasks require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
-Prior work established that standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., [1, 5, 6]), but there is no widely accepted approach that can generalize to unknown tasks [7].
-Generalizing across tasks is difficult primarily due to a shift in the relative balance of workload components; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1].
-This dynamic relationship between physiological signals and an individual’s underlying workload is exemplar of the variability of human behavior, which makes the development of machine learning models that generalize in the real world difficult.
-Gathering the necessary data to develop these models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of human activity within real world domains.
-Similar complexity is seen across human-robot interaction (HRI) application domains [8, 9, 10].
-The real world variability of human behavior and activity violates the IID assumption, requiring non-standard machine learning techniques (e.g., continual learning, meta-learning) to successfully overcome these challenges [11].
+Urban firefighters will work alongside robots to perform many tasks (Delmerico et al., 2019).
+Firefighters have historically used robots to gather aerial imagery, but advanced control algorithms will enable robots to extinguish fires in otherwise inaccessible locations (Perez-Saura et al., 2023).
+This new task will require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
+The dynamic relationship between physiological signals and an individual’s underlying workload makes the development of machine learning models that generalize in the real world difficult (Albuquerque et al., 2022; Popivanov and Mineva, 1999).
+Prior work demonstrated that this dynamism is considerably influenced by contextual factors (e.g., tasks demands, teaming dynamics), as contextual factors heavily influence an individual’s workload (Heard et al., 2018; Longo et al., 2022) and their physiological response to external stimuli (Larradet et al., 2020; Saganowski et al., 2022).
+Considering how contextual factors influence each workload components adds additional complexity.
+Each component relies on a unique subset of physiological signals (Bhagat Smith et al., 2022, 2023; Fortune et al., 2020; Heard et al., 2019), but it is impossible to predict the specific physiological response novel tasks will incur.
+Gathering the necessary data to develop workload estimation models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of real-world human activity (Bhagat Smith et al., 2024).
+The variability between contexts (e.g., tasks) violates the IID assumption (Albuquerque et al., 2022; Cao, 2014; Popivanov and Mineva, 1999), and must be analyzed through the lens of distribution shifts (i.e., the mathematical differences between datasets).
+Prior work developed a broad range of non-IID machine learning techniques (e.g., continual learning, meta-learning) that incorporate contextual examples to ameliorate a distribution shift’s negative impact (Quiñonero-Candela et al., 2022).
+These techniques have demonstrated success in many application domains (e.g., computer vision (Tian et al., 2020), natural language (Latif et al., 2022), task recognition (Leite and Xiao, 2022)), but their success varies based upon several dataset characteristics (e.g., labeling, volume).
+There exists too many techniques for a comprehensive evaluation to be tractable; thus, a more careful analysis evaluating how relevant human factors interact with a technique’s machine learning characteristics is required.
 The manuscript surveys applicable non-IID machine learning methods.
-Specifically, the non-IID methods are evaluated to assess their viability for developing models of the human’s internal state (e.g., workload estimation) that can accommodate real world variability.
-Workload estimation for unknown tasks is an exemplar of the challenge in constructing machine learning models for HRI application domains that are characterized by non-IID data.
-Section II discusses the fundamental of objective workload estimation and distribution shifts that characterize non-IID machine learning.
-Section III lays out the notation and evaluation criteria.
-Section IV evaluates a wide range of non-IID machine learning methods, and Section V discusses the most applicable methods for workload estimation for unknown tasks.
-Finally, Section VI concludes and presents potential future work.
-II Background
+These methods are evaluated to assess their viability for developing workload estimation models that can accommodate real-world variability.
+First, the fundamentals of objective workload estimation and distribution shifts are presented, followed by requisite notation and criteria.
+These criteria are used to analyze a wide range of non-IID machine learning techniques, where arguments for which techniques merit empirical investigation are presented.
+The most applicable techniques are further discussed, which highlights the importance of jointly evaluating a problem’s machine learning and human factors considerations in order to determine a techniques real-world applicability.
+Finally, conclusions and future work are presented.
+2 Background
 Workload is a complex, dynamic, individual-specific, non-linear construct, and developing models that can account for these complexities is a non-trivial task.
-Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown [7] [12].
-Modern techniques (e.g., meta-learning) approach machine learning from a different perspective and possess the potential to build more flexible workload estimation models.
+Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown (Longo et al., 2022; Zhao et al., 2018).
 Humans have limited resources for processing information, making decisions, and dealing with physical stress.
-Formally, workload is the degree of activation of a finite pool of resources while performing a task [7] that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience).
-Workload has been shown to vary across individuals, across tasks, and over time [13] [14].
-Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical [15], and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile [16].
-Workload estimation methods either rely on subjective (e.g., NASA Task Load Index [17], In-Situ Surveys [18]) or objective metrics.
-Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports [19] [20].
+Formally, workload is the degree of activation of a finite pool of resources while performing a task (Longo et al., 2022) that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience).
+Workload has been shown to vary across individuals, across tasks, and over time (Christensen et al., 2012; Wickens et al., 2004).
+Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical (Mitchell, 2000), and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile (Heard et al., 2018).
 Robots operating in dynamic environments need high-frequency workload measures to adequately adjust to their human teammates’ current state.
-Objective estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., [21, 22, 23, 24, 25]).
-Human-robot teams that are deployed in unstructured, dynamic environments cannot rely on environmentally embedded sensors (e.g., built-in cameras), or complex sensor systems (e.g., electroencephalogram (EEG)).
-Thus, wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently.
+Workload estimation methods either rely on subjective or objective metrics.
+Surveys and questionnaires (e.g., NASA Task Load Index (Hart and Staveland, 1988), In-Situ Surveys (Wilson and Russell, 2003)) are common subjective metrics that rely on an individual to self-assess their perceived workload levels (Heard et al., 2018).
+Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports (Kosch et al., 2023; Matthews et al., 2020); thus, they are not viable for real-world environments.
+Physiological signals provide a quantitative and consistent method for measuring a human’s workload and are the most widely used objective metric (Debie et al., 2019; Guan et al., 2022).
+Objective workload estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., Cao et al., 2021; Guo et al., 2021; Quan et al., 2023).
+These methods provide workload estimates at the necessary frequency and accuracy for real-world environments (Bagheri and Power, 2022; Fortune et al., 2020; Heard et al., 2019).
+Many objective workload estimation methods are ill suited for uncertain, dynamic environments as the sensors required to collect the necessary physiological signals are either complex systems (e.g., electroencephalogram (EEG) (Ved and Yildirim, 2021)) or environmentally embedded (e.g., built-in cameras (Kosmopoulos et al., 2012)).
+The human must be able to move through the environment freely.
+Environmentally embedded sensors restrict the human’s movement and complex sensor systems are extremely sensitive to the noise introduced by human movement, making them ill suited for uncertain, dynamic environments.
+Wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently.
+Wearable sensors are susceptible to high sensor noise and latency issues, but prior work has demonstrated their ability in accurately estimating workload in real-time environments (Dayal et al., 2024; Heard et al., 2020; Liu et al., 2024; Park et al., 2024).
+Robots need a reliable estimate of the distribution of workload across the workload components, as tasks will change over time, and differing tasks impact different components.
 Each workload component corresponds to a different set of physiological metrics, though some metrics may correspond to multiple components.
-A good example is pupil diameter, which is positively correlated with cognitive and visual workload [16].
-Robots need a reliable estimate of the distribution of workload across the components, as tasks will change over time, and differing tasks impact different components.
-Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability [16].
+Incorporating multiple wearable sensors and physiological signals allows a workload estimation method to differentiate between the workload components and accurately measure their contribution to overall workload (Heard et al., 2018; Bhagat Smith et al., 2024).
+Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability (Heard et al., 2018).
 Sensitivity represents an algorithm’s ability to reliably detect workload levels (i.e., $\geq$ 80% accuracy or $\leq$ 5 root mean squared error (RMSE)).
@@
 The last metric, generalizability, refers to the algorithm’s ability to generalize across individuals and tasks.
-This metric defines “generalizing across tasks” as an algorithms capacity to asses the complete overall workload state (i.e., each workload component) and defines “generalizing across individuals” as an algorithms capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
-Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., [26, 27, 28, 29]).
-Many methods are either suitable (e.g., [30, 31, 32, 33]) or sensitive (e.g., [34, 35, 5, 36]), but very few algorithms are diagnostic.
-Further, little work has been done to evaluate the generalizability of existing workload estimation methods.
-Prior work tends to focus on either generalizing across individuals or generalizing across tasks.
-Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., [37, 24, 38, 39]); however, generalizing across tasks, either known or unknown, is rarely evaluated.
-Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms [40].
-Much of this work evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., [40, 41, 42, 43, 44]).
-Thus, these algorithms are sensitive, but not generalizable.
-Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., [45, 46, 47]); thus, failing to meet the generalizability criterion.
-A support vector machine-based workload estimation algorithm that uses Recursive Feature Elimination to optimize the input features was shown to estimate cognitive workload on arbitrary tasks [48].
-This method achieved $<$ 1.0 mean squared error on cross-task workload estimation.
-Similar methods were developed by Guan et al.
-[23] and Zhou et al.
-[49].
-Additionally, deep learning techniques (e.g., attention mechanisms) that estimate workload have shown promise [50].
-However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from a static sensor (i.e., EEG [51, 52, 53]).
+This metric defines “generalizing across tasks” as an algorithm’s capacity to asses all seven workload components and defines “generalizing across individuals” as an algorithm’s capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
+Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., Dell’Agnola et al., 2021; Hogervorst et al., 2014; Xie et al., 2019; Yin et al., 2019).
+Many methods are either exhibit suitability (e.g., Islam et al., 2020; Momeni et al., 2019; Moustafa et al., 2017; Zhang et al., 2020) or sensitivity (e.g., Caywood et al., 2017; Ding et al., 2020; Kaczorowska et al., 2021; Ved and Yildirim, 2021), but very few exhibit diagnosticity.
+Further, little work evaluated the generalizability of existing workload estimation methods.
+Prior work tends to focus on generalizing across individuals.
+Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., Albuquerque et al., 2019; Guo et al., 2021; Heard et al., 2019; Hefron et al., 2017; Novak et al., 2015); however, generalizing across tasks, either known or unknown, is rarely evaluated.
+Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms (Appel et al., 2023).
+Many of these efforts evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., Appel et al., 2023; Baldwin and Penaranda, 2012; Besson et al., 2012; Boring et al., 2020; Walter et al., 2013).
+Thus, these algorithms exhibit sensitivity, but not generalizability.
+Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., Dimitrakopoulos et al., 2017; Kakkos et al., 2021; Zhang et al., 2018); thus, failing to meet the generalizability criterion.
+Some support vector machine-based workload estimation algorithms have achieved achieved $<$ 1.0 mean squared error on cross-task workload estimation (Guan et al., 2022; Ke et al., 2015; Zhou et al., 2022).
+Deep learning techniques (e.g., attention mechanisms) that estimate workload have also shown promise (Taori et al., 2022).
+However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from complex or environmentally embedded sensors (e.g., Ji et al., 2023; Guan et al., 2023; Wen et al., 2023).
 Thus, none of the methods are viable for use in dynamic, uncertain environments.
-Real world human-robot teams will inevitably encounter unique scenarios when deployed.
-An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences [14].
-Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples [2].
-Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1].
-This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques [2].
-III Notation and Evaluation Criteria
-Dynamic human-robot teams will rely on machine learning methods to enable a robot to quantify different aspects of their human teammate (e.g., workload).
-A major challenge is that these teams will continuously encounter new scenarios that the robot’s underlying machine learning methods cannot account for.
-These new scenarios constitute a meaningful distribution shift from conventional machine learning method’s training dataset (i.e., violate the IID assumption).
-Specifically, a distribution shift indicates scenarios where training and testing data come from different distributions.
-Understanding the difference between non-IID machine learning methods is key for building human-robot teams that can handle the variability of real world problem domains.
+Real-world human-robot teams will inevitably encounter unique tasks when deployed.
+An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences (Wickens et al., 2004).
+Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples (Murphy, 2023, chap.
+19).
+Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations (Heard et al., 2019).
+This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques (Murphy, 2023, chap.
+19).
+3 Notation and Evaluation Criteria
+Dynamic human-robot teams will employ robots that estimate their human teammate’s workload using machine learning methods.
+Extending the robot’s capabilities to accommodate unknown tasks poses unique machine learning challenges, as these tasks constitute a meaningful distribution shift.
+A distribution shift indicates a scenario where training and testing data come from different probability distributions.
+Understanding the difference between applicable machine learning methods is key for building human-robot teams that can handle the real-world complexity.
 The workload estimation and the machine learning literature both cover a large spectrum of work.
-Naturally, there are terms that frequently occur in both fields which require disambiguation.
-The term task is used in the workload literature to refer to a job an individual may perform (e.g., lifting boxes, mental calculations), but the machine learning literature uses the term to refer to a learning task for a machine learning model.
-The term domain is often used to characterize the input space for a machine learning problem.
-For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather [54].
-Domain may also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
-The term learning task will refer to instances of a machine learning problem.
-The term task, without any adjectives, will refer to a task conducted by a human.
+There are frequently occurring terms in both fields; thus, analyzing applicable methods requires disambiguating key terminology and establishing evaluation criteria to ensure consistent analysis.
+3.1 Notation
+The term task in the workload literature refers to a job an individual may perform (e.g., lifting boxes), but the machine learning literature uses the term to refer to a learning task for a machine learning model.
+The term domain is used to characterize the input space for a machine learning problem.
+For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather (Sun et al., 2022).
+Domain can also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
+The term learning task will refer to a machine learning problem.
+The term task, without adjectives, will refer to an activity conducted by a human.
 The term domain will refer to different input spaces that a machine learning algorithm may operate on.
 The term application domain will refer to applications where human-robot teams may be deployed.
-Definitions of domain and learning task, in the context of distribution shift, have long been established [55]:
-Definition III.1
-(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},…,x_{n}\}\in\mathcal{X}$.
-Definition III.2
+The definitions of domain and learning task, in the context of distribution shift, are established (Pan and Yang, 2009):
+Definition:.
+(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},\dots,x_{n}\}\in\mathcal{X}$.
+Definition:.
 (Learning Task) A learning task $L$ is a two-tuple $(\mathcal{Y},f(x))$ associated with a specific domain $D$.
@@
 $f(x)$ can be written as a conditional probability distribution $P(y|x)$.
-Grounding this notation in workload estimation helps to provide an intuitive explanation of the different variables.
+Grounding this notation in workload estimation provides an explanation of the different variables.
 A particular workload estimation instance represents the domain, $D$, which is characterized by the set of physiological metrics, $\mathcal{X}$.
@@
 Changes to the predictive function $f(x)$ are less intuitive and more easily explained via task recognition.
-A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., (1, 2, and 3)), where each integer represents a unique task.
-A similar task recognition problem can be defined using the same metrics and the same label space, but where each integer represents the task difficulty (i.e., easy, medium, or hard).
-These two task recognition problems are semantically different, but the underlying machine learning method does not inherently recognizes these differences; therefore, uniqueness of these two learning tasks cannot be fully specified by the domain and label space.
+A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., [1, 2, and 3]), where each integer represents a unique task.
+A similar task recognition problem can be use the same metrics and the same label space, but with each integer representing the task difficulty (i.e., [easy, medium, or hard]).
+These two problems are semantically different, but the underlying machine learning method does not inherently recognize these differences; therefore, the uniqueness of these two learning tasks cannot be fully specified by the domain and label space.
 The changes in how the resulting predictive functions, $f(x)$, map the physiological metrics to the output values is a key characteristic in differentiating these two learning tasks.
-Numerous methods have been applied to the non-IID machine learning problems, each characterized by the number of distributions and distribution shift type (see Fig.
-1).
-Three evaluation criteria are applied to the presented methods: (1) Portability (2) Model Complexity (3) Adaptability.
-Portability refers to the number of distributions used to train and test a model [56].
-Learning tasks may consist of multiple source domains (e.g., rural hospital data vs.
-urban hospital data).
+3.2 Evaluation Criteria
+An exhaustive analysis of applicable machine learning methods is impractical due to the volume of existing methods applied to non-IID machine learning problems; however, these methods can be conceptually grouped based upon the number of distributions and the distribution shift type (see Fig.
+1) (Murphy, 2023).
+These categories represent a comprehensive list of non-IID machine learning method types and each category is assessed to understand which method types merit further investigation.
+Three evaluation criteria are applied to the presented techniques: (1) Portability, (2) Model Complexity, and (3) Adaptability.
+These three criteria function as a tool for understanding how the human factors considerations of workload estimation for unknown tasks interact with the machine learning aspects of non-IID techniques, with the aim of determining each technique’s practical utility.
+Portability refers to the number of distributions used to train and test a model (Sun et al., 2015).
+Learning tasks may consist of multiple source domains (e.g., urban vs wildland fire fighting).
 A straightforward means of combining information from multiple domains is to construct a single joint distribution; however, data in some domains may be of higher quality than others.
-Weighting high-quality data over low-quality data is appropriate to maximize performance [56].
+Weighting high-quality data over low-quality data is appropriate to maximize performance (Sun et al., 2015).
 Additionally, the separation of source domains may be artificial.
-Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks [57].
+Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks (Baxter, 2000).
 Portability can have four values: one-to-one, one-to-many, many-to-one, and many-to-many.
@@
 Many-to-many knowledge transfer occurs between multiple sources to multiple target domains simultaneously.
-Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of [2].
-Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful [58].
+A many-to-many portability is desirable for workload estimation for unknown tasks.
+Machine learning-based workload estimation methods are primarily developed using data collected in experimental conditions, producing diverse datasets with low data volumes.
+Further, these experimental conditions may unintentionally introduce bias into the dataset (Bhagat Smith et al., 2024).
+Developing workload estimation methods must therefore be able to accommodate distribution shifts for multiple, small, noisy datasets (i.e., many-to-many portability).
+Many-to-one portability is only desirable if a method needs to estimate workload for a single new task, which is unlikely in real-world application domains.
+One-to-one portability is not sufficient outside of contrived circumstances, as effectively capturing the nuanced differences between known and unknown tasks using two distributions requires large data volumes.
+Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of (Murphy, 2023, chap. 19).
+Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful (Goodfellow et al., 2016).
 Machine learning models with more parameters require higher volumes of training data.
-Early statisticians built simple linear models using hundreds of examples [58], whereas modern deep learning architectures may require tens of billions of examples [59].
-Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as more complex models are often required to solve more complex problems.
-Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for machine learning model development; thus, smaller models are preferred.
-An example of low model complexity is a random forest model that contains a few hundred decision trees, where the depth of each tree is dataset-dependent; thus, these models typically contain on the order of thousands of parameters [60].
-An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters [22], but may contain on the order of millions of parameters.
-An example of high model complexity is deep neural networks (e.g., convolutional networks, long short-term memory networks (LSTM)), also referred to as deep learning architectures.
-Deep learning architectures can contain anywhere from tens of millions [61] to over 100 billion parameters [59].
-The last metric, Adaptability, refers to the amount of data required to update a machine learning model to solve new problems [62] [63].
-Many non-IID machine learning methods seek to account for distribution shifts by updating existing machine learning models’ parameters so that the model can solve new learning tasks.
+Additionally, training machine learning models in human-robot interaction (HRI) application domains requires this data to be ecologically valid, introducing additional complexity to model development.
+Early statisticians built simple linear models using hundreds of examples (Goodfellow et al., 2016), whereas modern deep learning architectures may require tens of billions of examples (Brown et al., 2020).
+Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as complex models are often required to solve complex problems.
+Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for data collection; thus, smaller models are preferred.
+An example of low model complexity is a random forest model that contains a few hundred decision trees and typically contain on the order of thousands of parameters (Breiman, 2001).
+An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters (Fortune et al., 2020), but may contain on the order of millions of parameters.
+Two examples of high model complexity are deep learning and generative AI architectures (e.g., convolutional networks, long short-term memory networks (LSTM), transformers).
+These architectures can contain anywhere from tens of millions (Wilson and Cook, 2020) to over 100 billion parameters (Brown et al., 2020).
+A machine learning-based workload estimation method must be capable of running onboard a robot to be useful in uncertain, dynamic environments.
+The robot is not guaranteed to have a remote connection to a powerful computer or access to the cloud.
+Modern edge compute solutions (e.g., NVIDIA Jetson Orin (NVIDIA, 2025)) have the capability of executing many deep learning architectures in real-time, but it is desirable to minimize the model complexity in order to reduce the requisite computational resources and eliminate reliance on cloud resources.
+Further, workload estimation methods will be incorporated into a larger human-robot teaming architecture that may rely on other machine learning models.
+Reducing the model complexity of the workload estimation method frees up resources required by other components.
+Minimizing model complexity reduces the power consumed by the machine learning methods allow for more efficient power consumption, a critical constraint in uncertain, dynamic environments.
+Smaller models also require fewer computational resources that may be necessary for other learning tasks (e.g., task recognition, decision making).
+The last metric, Adaptability, refers to the amount of data required to update a machine learning model for new learning tasks (Wang et al., 2020b).
+Many non-IID methods seek to account for distribution shifts by updating existing machine learning models’ parameters.
 Adaptability is the number of test data points necessary for the updated model to make accurate predictions for new learning tasks, where the goal is to minimize the required data.
 A machine learning method has no adaptability if it requires the same amount of training data in the source and target domains.
-Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance [58].
-Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain [63].
+Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance (Goodfellow et al., 2016).
 A 50% reduction in data required for updating a model is non-trivial, and may still constitute a large dataset.
-An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts [62].
-It is important to note adaptability is application domain-specific, as the learning tasks are fundamentally different.
-IV Non-IID Machine Learning Methods
-There are several non-IID machine learning methods that can account for different distribution shift types, including: transfer learning, domain adaptation, test-time training, continual learning, domain generalization, and few-shot learning.
-Each method corresponds to a specific distribution shift type and the number of distributions considered (see Figure 1).
-There are two primary distribution shifts: (1) Domain shift, and (2) Label shift.
-A domain shift indicates a distribution shift of the input features, and label shift indicates a shift of the output values.
+Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain (Zhuang et al., 2020).
+An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts (Wang et al., 2020b).
+Adaptability is application-domain specific, as learning tasks are fundamentally different.
+Real-world human-robot teams will encounter unknown tasks in the field and must be able to estimate the human’s workload as quickly as possible.
+Achieving a high adaptability is ideal, requiring a workload estimation method to collect large amounts of data in dynamic environments is impractical.
+4 Non-IID Machine Learning Methods
+Non-IID machine learning methods seek to account for distribution shifts by intelligently incorporating information from multiple distributions (i.e., domains) and fall into six categories: domain adaptation, transfer learning, test-time training, continual learning, domain generalization, and few-shot learning (see Figure 1).
+These categories are primarily differentiated by their data handling procedure, which can also be viewed as how each method formulates the distribution shift problem (Murphy, 2023, chap.
+19).
+Analyzing each category’s data handling procedures is central to understanding a method’s practical utility for estimating workload associated with unknown tasks, as it provides insight into how these data handling procedures can take advantage of the unique characteristics of HRI datasets.
+Datasets used to train machine learning-based workload estimation methods are heavily impacted by the variability inherent in the physiological signals, noise within wearable sensor data, and range of potential human behaviors (Das Chakladar and Roy, 2024).
+These data characteristics present challenges unique to HRI application domains and differentiate them from other learning tasks (e.g., object detection, speech recognition).
+Additionally, most machine learning-based workload estimation methods rely on datasets collected in controlled experimental conditions.
+Methods within each category implement similar data handling procedures, where each category varies based upon the distribution shift type and number of distributions considered.
+There are two primary distribution shift types.
+A domain shift indicates a distribution shift of the input features (e.g., heart-rate variability, pupil diameter), and label shift (e.g., workload value) indicates a shift of the output values.
+Domain shifts can be characterized by changes to the range of values in various physiological signals or changes to the relative importance for any given physiological signal for estimating workload.
 Label shifts can be characterized by either a new distribution of output values, or a new set of values.
-Standard IID methods only account for data from a single distribution, which can lead to performance degradation in the presence of distribution shifts [2].
-Non-IID methods seek to account for these shifts by intelligently incorporating information from multiple distributions (i.e., domains).
-Further, these methods either account for exactly two distributions or from more than two distributions, where the boundaries between distributions may be real or artificial.
-The variability of human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments.
+For example, consider a training dataset where the cognitive workload values range from 0 to 15, with a mean value of 10 and a median value of 5.
+A testing dataset with a label shift may have cognitive workload values ranging from 5 to 20, or may have a mean and median value of 14.
+The degree to which these distribution shifts impact performance varies based upon the shift’s magnitude and dataset size.
+Non-IID machine learning methods either account for exactly two distributions or more than two distributions, where the boundaries between distributions may be real or artificial.
+Human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments.
 However, the data can be decomposed or re-organized such that any of the presented techniques can be applied.
-Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks [64], or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models [63].
-Many of these methods have been successfully applied to task recognition using wearable sensors (e.g., [65, 66, 67, 68]), which shares core commonalities with workload estimation.
-Thus, these methods provide insight into the capabilities of the underlying machine learning models to account for distribution shifts in relevant application domains.
-Prior work has primarily focused on evaluating standard IID methods’ ability to generalize to new situations, where many methods fail to generalize to new tasks (e.g., [40, 43, 48, 47]).
-Relying on standard IID methods ignores the distribution shifts of the physiological signals and workload components as tasks change.
-No prior work modeled workload estimation for unknown tasks as a non-IID problem.
-An appropriate non-IID machine learning method for the workload estimation for unknown tasks must be identified, so that the resulting workload estimation algorithm can account for distribution shifts across tasks.
-IV-A Domain Adaptation
-Domain adaptation considers the setting where training data and test data are characterized by the same features, but are sampled from different distributions.
-Definition IV.1
-(Domain Adaptation [69]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$.
+Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks (Vinyals et al., 2016), or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models (Zhuang et al., 2020).
+An analysis of non-IID machine learning methods that considers the unique characteristics of HRI datasets and how they interact with each method’s data handling procedure is presented.
+Relevant methods within each category are briefly discussed to assess their merit for estimating human workload for unknown tasks using the developed criteria (i.e., portability, model complexity, and adaptability).
+4.1 Domain Adaptation
+Domain adaptation considers scenarios where training and testing data contain the same features (e.g., physiological signals), but are sampled from different distributions.
+Definition:.
+(Domain Adaptation) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$ (Singhal et al., 2023).
 This distribution shift type is called a domain shift, or a covariate shift.
-For example, consider multiple datasets with images of the same objects [70].
-Changes in the images’ lighting, background, or orientation represent meaningful difference in the input features and are prototypical examples of a domain shift.
+An intuitive example is a machine learning-based workload estimation method trained on data collected in an climate controlled environment (i.e., source domain) and tested on data in an outdoor environment (i.e., target domain).
+Extreme temperature differences will have an impact on an individual’s workload, but also alter their physiological responses to nominal workload conditions (Hancock and Vasmatzidis, 2003).
 Domain adaptation is a non-IID machine learning method that seeks to minimize performance degradation in the presence of domain shift.
 It is important to note that domain adaptation techniques focus on distribution shifts between one source and one target domain (i.e., two distributions), as shown in Figure 1.
-Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training in order to perform well [55].
-Further, there are two forms of domain adaptation: supervised [70] and unsupervised [61].
-Supervised domain requires that labels be available for $L_{T}$, and there are no labels available for $L_{T}$ for unsupervised domain adaptation.
-There tends to be significantly more data available in source domains than the target domains, and using labeled data from both domains typically results in overfitting to the source distribution [70].
+Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training to perform well (Pan and Yang, 2009).
+Further, there are two forms of domain adaptation: supervised (Wang and Deng, 2018) and unsupervised (Wilson and Cook, 2020).
+Supervised domain adaptation requires that labels be available for $L_{T}$, while there are no labels available for $L_{T}$ for unsupervised domain adaptation.
+There tends to be significantly more data available in source domains, and using labeled data from both domains typically results in overfitting to the source distribution (Wang and Deng, 2018).
 Thus, the majority of prior work has focused on unsupervised domain adaptation.
-Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., [71, 72, 65, 73, 74]), workload estimation (e.g., [49, 75, 76]), and emotion recognition (e.g., [77, 78, 79, 80]).
-The vast majority of these applications use adversarial learning techniques [66, 75, 80], though self-supervised and feature representation learning strategies have also been moderately successful [77].
-Adversarial domain adaptation consists of three model types: (1) feature extractor, (2) domain discriminator, and (3) label predictor.
-Seminal work integrated adversarial learning and domain adaptation as a mini-max game [81].
-This work demonstrated that domain adaptation can be achieved by jointly optimizing a feature extraction model to maximize the difference of latent features from source and target domain data, via a domain classifier, and to minimize the error in predicted values for learning tasks, via a label predictor.
-This joint optimization enables the effective learning of a useful latent feature representation.
-Prior work demonstrated that replacing the discriminative feature extraction models with generative feature extraction models is also an effective means of domain adaptation (e.g., [82, 83, 84]).
-Most domain adaptation techniques exhibit one-to-one portability [84], but multi-target domain adaptation (i.e., one-to-many portability) is an active research area [85].
-Domain adaptation techniques primarily rely on deep neural networks [49, 75, 77], consisting of millions of parameters; thus, exhibiting a high model complexity.
-The adaptability of domain adaptation varies.
-Some approaches require one example in the target domain [86], while others require thousands [66].
-Further, access to source and target data during training is a hard requirement for many domain adaptation techniques, which makes them impractical for real world HRI.
+Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., Alajaji et al., 2023; An et al., 2021; Faridee et al., 2022; Zakia and Menon, 2021; Li et al., 2022b), workload estimation (e.g., Zhou et al., 2022, 2023b; Albuquerque et al., 2022), and emotion recognition (e.g., Li et al., 2022c; Latif et al., 2022; Guo et al., 2023; He et al., 2022).
+The vast majority of these applications use adversarial learning techniques (Li et al., 2021c; Zhou et al., 2023b; He et al., 2022), though self-supervised and feature representation learning strategies have also been moderately successful (Li et al., 2022c).
+Most domain adaptation techniques exhibit one-to-one portability (Balaji et al., 2020; Liu and Tuzel, 2016; Li et al., 2020), but multi-target domain adaptation (i.e., one-to-many portability) is an active research area (Gholami et al., 2020).
+Domain adaptation techniques primarily rely on deep neural networks (Zhou et al., 2022, 2023b; Li et al., 2022c) consisting of millions of parameters; thus, exhibiting a high model complexity.
+Domain adaptation techniques exhibit varying adaptability, where some approaches require one example in the target domain (Yue et al., 2021), while others require thousands (Li et al., 2021c).
+Further, access to source and target data during training is a core requirement for all domain adaptation techniques, making them impractical for real-world HRI.
 One-to-one portability, high model complexity, and low adaptability are undesirable properties; thus, domain adaptation techniques are not viable for workload estimation of unknown tasks.
-IV-B Transfer Learning
+4.2 Transfer Learning
 Transfer learning aims to improve the performance of a machine learning model for a target domain by transferring knowledge from a different, but related source domain.
-Definition IV.2
-(Transfer Learning [63]) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$.
-Transfer learning is characterized a label shift [87], which can refer to a shift in the relative distribution of output values (e.g., disease prevalence in a rural hospital vs.
-urban hospital), or a complete change in the labels available.
-Transfer learning techniques are capable of transferring knowledge from a model trained to identify everyday objects (e.g., cats, planes) in images to a model performing medical image analysis (e.g., ultrasounds, x-rays) [88].
-Similar to domain adaptation, transfer learning has been widely applied to HRI (e.g., human task recognition [89, 90], emotion recognition [91, 92, 93, 94]).
-There are many transfer learning techniques, which can be categorized based on how knowledge is transferred: instance-based (e.g., [95, 96, 97]), feature-based (e.g., [98, 99, 100]), and parameter-based (e.g., [101, 87, 68, 102]).
-Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques [103].
+Definition:.
+(Transfer Learning) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$ (Zhuang et al., 2020).
+Transfer learning is characterized by a label shift (Huh et al., 2016), either in the relative distribution of output values or a complete change in the available labels.
+Consider the field of affective computing’s use of physiological signals (Egger et al., 2019).
+Transfer learning techniques attempt to transfer knowledge from a model trained to recognize positive emotions (e.g., awe) to a model attempting to recognize negative emotions (e.g., anger, disgust).
+These techniques are widely applied within HRI application domains (e.g., human task recognition (Ding et al., 2018; Li et al., 2021b), emotion recognition (Li et al., 2019a; Quan et al., 2023; Ma et al., 2023; Nguyen et al., 2023)).
+There are many transfer learning techniques that vary based on how knowledge is transferred: instance-based (e.g., Chen et al., 2022c; Wang et al., 2019a, b), feature-based (e.g., Ren et al., 2019; Stüber et al., 2018; Wu et al., 2021), and parameter-based (e.g., Fu et al., 2021; Huh et al., 2016; Soleimani and Nazerfard, 2021; Link et al., 2022).
+Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques (Weiss et al., 2016).
 Parameter-based methods have two phases: pre-training and fine-tuning.
-Pre-training trains a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$.
+Pre-training optimizes a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$.
 Fine-tuning trains the same model on data from the target domain, $D_{T}$, but only updates a subset of the parameters.
 This two-step training procedure enables the network to learn a useful latent representation during pre-training, and a task-specific transformation of that representation during fine-tuning.
-This approach has been successful in computer vision [87], robotics [104], and task recognition (e.g., [105, 106, 107, 108]).
+This approach has been successful in computer vision (Huh et al., 2016), robotics (Lee et al., 2021), and task recognition (e.g., An et al., 2023; Ray et al., 2023; Pavliuk et al., 2023).
 Many of these methods use deep learning architectures to transfer knowledge across source/target domain pairs, requiring thousands of examples to be successful.
-Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability [63] [59].
+Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability (Zhuang et al., 2020; Brown et al., 2020).
 Zhuang et al.
-evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability [63].
-Similar findings exist in transfer learning for human activity recognition [105].
-Similar techniques were also developed to solve a wide range of natural language problems based on only a few examples [59].
-This high adaptability is in direct contrast to the previously discussed transfer learning techniques, where the primary difference is the model complexity.
-A transformer deep learning architecture was trained to predict the next word in a sentence [59].
+evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability (Zhuang et al., 2020).
+Similar techniques were developed to solve a wide range of natural language problems based on only a few examples (Brown et al., 2020).
+This high adaptability is in direct contrast to the other transfer learning techniques, where the primary difference is the model complexity.
+A transformer deep learning architecture was trained to predict the next word in a sentence (Brown et al., 2020).
 This network was the foundation of several downstream networks that performed tasks, including machine translation, question answering, and arithmetic word problem-solving.
-The architecture consisted of 175 billion parameters, requiring significantly more data to train parameter-based techniques that rely on smaller networks (i.e., tens of billions of data points).
+The architecture consisted of 175 billion parameters, requiring significantly more data to train than the parameter-based techniques that use smaller networks (i.e., tens of billions of data points).
 These results suggest that a higher model complexity can directly improve portability and adaptability.
@@
 It is also desirable that workload estimation algorithms exhibit many-to-many portability, as models may need information across multiple individuals and tasks.
-Thus, transfer learning is not a viable option for workload estimation of unknown tasks.
-IV-C Test-time Adaptation
+Thus, transfer learning is not a viable option.
+4.3 Test-time Adaptation
 Domain adaptation techniques learn to anticipate domain shifts by simultaneously training on $D_{S}$ and $D_{T}$; however, target domain data may not be available during training.
 Test-time adaptation techniques seek to overcome this limitation.
-Definition IV.3
-(Test-time Adaptation [109]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$.
+Definition:.
+(Test-time Adaptation) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$ (Sun et al., 2020).
 Test-time adaptation techniques seek to overcome domain shifts by separating the training process into two distinct phases: (1) training on $D_{S}$ data, and (2) updating that model with $D_{T}$ data.
-This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 [109].
-Prior work applied test-time adaptation techniques to egocentric task recognition [110] and adaptive policy optimization for assistive robots [111].
+This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 (Sun et al., 2020).
+Prior work applied test-time adaptation techniques to egocentric task recognition (Plananamente et al., 2022) and adaptive policy optimization for assistive robots (He et al., 2023).
 However, test-time adaptation techniques have not been widely applied to HRI application domains.
@@
 These techniques can be categorized into as either self-supervised, or entropy-based.
-Self-supervised techniques use a dual-output neural network and a two-phase training procedure [109].
+Enumerating the algorithmic details for each of these techniques will highlight how they may be used in future human workload estimation algorithms.
+Self-supervised techniques use a dual-output neural network and a two-phase training procedure (Sun et al., 2020).
 Labeled data from the $D_{S}$ undergoes some deterministic transformation (e.g., image rotation).
@@
 This two-phase procedure decouples training from adaptation and enables the model to adapt at test-time.
-Many techniques take inspiration from this work (e.g., [112, 113, 114]); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
-Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., [115, 116, 117]).
-Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy [118] is calculated for each batch.
-The adaptation phase updates the shared feature extractor to minimize the average entropy over all batches [119].
+Many techniques take inspiration from this work (e.g., Azimi et al., 2022; Chen et al., 2023; Segu et al., 2023); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
+Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., Liu et al., 2021; Zhang et al., 2022, 2023a).
+Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy (Shannon, 1948) is calculated for each batch.
+The adaptation phase updates the shared feature extractor to minimize the mean entropy over all batches (Wang et al., 2020a).
+Other techniques use the same approach, but minimize Shannon Entropy from $D_{S}$ data and $D_{T}$ data (Jung et al., 2023; Seto et al., 2024).
+Prior work demonstrated that minimizing entropy in this way allows the feature extraction model to learn domain-invariant features to improve generalization to new learning tasks (Adachi et al., 2023).
+Recent work has extended entropy-based techniques to incorporate varying methods for comparing data points between $D_{S}$ and $D_{T}$ (e.g., Song et al., 2023; Niu et al., 2022; Zhang et al., 2023b).
 Fundamentally, test-time adaptation is domain adaptation without access to target domain data during training.
@@
 Prior work has primarily focused on deep learning architectures, but these techniques are general enough to apply to shallow neural networks.
-It is possible that smaller networks may possess sufficient adaptability to account for real world variability, but the efficacy of test-time adaptation on these networks has not been thoroughly evaluated.
-IV-D Continual Learning
+It is possible that smaller networks may possess sufficient adaptability to account for real-world variability, but the efficacy of test-time adaptation on these networks has not been evaluated.
+4.4 Continual Learning
 Continual Learning techniques train a model to perform a sequence of non-IID learning tasks, where each task is characterized by a unique domain shift (see Figure 1).
-Definition IV.4
-(Continual learning [120]) Given a sequence of domains $D$ = $D_{s_{1}},...,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},…,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$.
-These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks [121].
-Continual learning is also referred to as lifelong learning, sequential learning, or online learning [122].
-There is a wide range of techniques, including knowledge distillation (e.g., [123, 124, 125]), memory-based methods (e.g., [126, 127]), and parameter isolation (e.g., [128, 129]).
-Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., [130, 131, 132, 133]), affective computing (e.g., [8, 9, 134]), and adaptive human-robot interactions (e.g., [8, 135, 136]).
-Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary [137].
-Thus, continual learning techniques can be trained offline as the number of tasks is known a priori.
-These techniques are referred to as multi-head techniques, as separate models are trained for each learning task [138].
-However, knowing all the tasks ahead of time is not always realistic.
-Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online [138].
-These techniques are referred to as single-head techniques, since a single model is trained for all tasks.
+Definition:.
+(Continual learning) Given a sequence of domains $D$ = $D_{s_{1}},\dots,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},\dots,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$ (Chen and Liu, 2018).
+These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks (Smith et al., 2023).
+Continual learning is also referred to as lifelong learning, sequential learning, or online learning (Mai et al., 2022).
+There is a wide range of techniques, including knowledge distillation (e.g., Chen et al., 2022a; Wu et al., 2019; Zhu et al., 2021), memory-based methods (e.g., Cai et al., 2021; Derakhshani et al., 2021), and parameter isolation (e.g., Lomonaco et al., 2021; Serra et al., 2018).
+Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., Hasan and Roy-Chowdhury, 2015; Ye et al., 2019; Ashry et al., 2020; Leite and Xiao, 2022), affective computing (e.g., Churamani et al., 2020; Gao et al., 2023; Churamani et al., 2022)), and adaptive HRI (e.g., Churamani et al., 2020; Lesort et al., 2020; Spaulding et al., 2021).
+Further, probabilistic continual learning techniques are common in the cognitive science and psychology communities to construct human decision making models (e.g., Austerweil and Griffiths, 2013; Collins and Frank, 2013; Liew et al., 2016; Kanwal et al., 2022; Navarro et al., 2006; Niv, 2019).
+Specifically, Dirichlet process mixture models that function as a non-parametric Bayesian clustering algorithm that can recognize existing classes and detect out-of-distribution classes (Li et al., 2019b).
+Recent work demonstrated that these techniques can be combined with deep-learning models to improve performance for computer vision tasks (e.g., Jerfel et al., 2019; Xu et al., 2024; Willes et al., 2022).
+Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary (De Lange et al., 2021).
+Thus, continual learning techniques can be trained offline, as the number of tasks is known a priori.
+These techniques are called multi-head techniques, as separate models are trained for each learning task (Chaudhry et al., 2018).
+However, knowing all the tasks a priori is not always realistic.
+Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online (Chaudhry et al., 2018).
+These single-head techniques train a single model is trained for all tasks.
 Learning tasks within a sequence can either be independent or dependent.
@@
 Dependent learning tasks represent a sub-sequence of learning tasks that must be executed in a particular order, where learning task $L_{i}$ holds key information about learning task $L_{i+1}$.
-Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques [139, 140].
-The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes [136], shallow neural networks [127], and deep networks [122]).
-Many techniques exhibit low adaptability, requiring thousands of examples to perform well [122].
+Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques (Mundt et al., 2023; Ao and Fayek, 2023).
+The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes (Spaulding et al., 2021), shallow neural networks (Derakhshani et al., 2021), and deep networks (Mai et al., 2022)).
+Many techniques exhibit low adaptability, requiring thousands of examples to perform well (Mai et al., 2022).
 Portability is challenging to define for continual learning, as it exhibits many-to-many portability, but does so only in a sequential fashion.
@@
 Portability and adaptability are key qualities required to estimating workload successfully for unknown tasks; thus, continual learning is a subpar option.
-Figure 2: Comparison of applying the technique Model Agnostic Meta-learning [141] to: (a) Domain Generalization, sub-problems experience a domain shift, and (b) Few-shot Learning, sub-problems experience a domain and label shift.
-IV-E Domain Generalization
-Domain generalization considers the setting where training data from multiple source domains is characterized by the same features that are sampled from different distributions [142].
-Specifically, domain generalization is the process of transferring knowledge acquired from several related source domains and applying it to previously unknown domains [143], as indicated in Figure 1.
-Definition IV.5
-(Domain Generalization [142]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$.
+4.5 Domain Generalization
+Domain generalization considers the setting where training data from multiple source domains is characterized by the same features, but sampled from different distributions (Li et al., 2018a).
+Specifically, domain generalization is the process of transferring knowledge from several related source domains and applying it to previously unknown domains (Muandet et al., 2013), shown in Figure 1.
+Definition:.
+(Domain Generalization) Given a set of source domains $DS$ = $D_{s_{1}},\dots,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},\dots,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ (Li et al., 2018a).
 The goal is to leverage unique cross-domain information to improve the machine learning model’s generalization.
-These methods assume the construction of a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains.
+These methods assume that constructing a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains.
 A key feature of domain generalization is the lack of labeled target domain samples.
-Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., [144, 145, 146]), data manipulation (e.g., [147, 148, 149]), and learning strategy (e.g., [150, 151, 152]).
-The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned [153].
-Domain alignment techniques measure this difference in many ways, including minimizing moments [154], contrastive loss [155], KL divergence [156], or maximum mean discrepancy [157].
-Prior work combined domain alignment metrics with domain-adversarial learning [157].
-The primary difference between domain-adversarial techniques for domain generalization and domain adaptation is that the source/target discriminator must differentiate between multiple source domains.
+This setting exhibits natural similarities to human workload estimation for unknown tasks, as novel tasks will incur a novel physiological response not encountered in the training dataset (i.e., domain shift).
+Additionally, it is unlikely that labeled data will be available when attempting to estimate workload for unknown tasks in real-world environments.
+Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., Suh et al., 2023; Li et al., 2022a; Gu et al., 2023b), data manipulation (e.g., Ilse et al., 2020; Li et al., 2021a; Yue et al., 2019), and learning strategy (e.g., Khandelwal and Yushkevich, 2020; Sicilia et al., 2023; Liu et al., 2020).
+The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned (Hu et al., 2020).
+Domain alignment techniques measure this difference in many ways, including minimizing moments (Ghifary et al., 2016), contrastive loss (Kim et al., 2021), or maximum mean discrepancy (Li et al., 2018b).
+A key difference between domain-adversarial techniques for domain generalization and domain adaptation is that the discriminator must differentiate between multiple source domains.
 Data manipulation techniques seek to artificially enhance the size and diversity of a dataset via data augmentation or generation.
-Data augmentation performs deterministic transformations over existing data to create new examples [158].
+Data augmentation performs deterministic transformations over existing data to create new examples (Otálora et al., 2019).
 However, data augmentation is not domain-agnostic, as useful transformations are specific to an application domain.
-Data generation uses generative machine learning models (e.g., Variational Autoencoders [147], Generative Adversarial Networks [148]) to create completely new training examples.
-These techniques are domain-agnostic, but the underlying models are hard to train and require a high training data volume [159].
+Data generation uses generative machine learning models (e.g., Variational Autoencoders (Ilse et al., 2020), Generative Adversarial Networks (Li et al., 2021a)) to create completely new training examples.
+These techniques are domain-agnostic, but the underlying models are hard to train and require a high data volumes (Goodfellow et al., 2020).
 These techniques cannot be applied to workload estimation, as generating realistic physiological metrics that correctly correspond to the workload values is non-trivial.
-Learning strategy-based domain generalization techniques seek to improve cross domain generalization through novel training methodologies, including: 1) Self-supervsied learning 2) Ensemble learning, and 3) Meta-learning.
-Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation (see Section III.C).
+Learning strategy-based domain generalization techniques seek to improve generalization through novel training methodologies, including: 1) Self-supervised learning 2) Ensemble learning, and 3) Meta-learning.
+Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation.
 Ensemble learning is the process of training multiple machine learning models with varying weight initializations or training data splits, then aggregating the models’ outputs to make predictions.
-These extensively studied techniques are notoriously computationally expensive [160] and are not viable solutions for real-time HRI domains.
-Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains [141].
-Broadly, meta-learning is ‘learning to learn’ [161].
-Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) [141]) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
+These techniques are notoriously computationally expensive (Zhou, 2012) and are not viable solutions for real-time HRI domains.
+Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains (Finn et al., 2017).
+Broadly, meta-learning is ‘learning to learn’ (Hospedales et al., 2021).
+Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) (Finn et al., 2017)) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
 A meta-model, parameterized by $\theta$, is trained to predict an initialization for the parameters of another model, $\phi_{i}$, that performs the learning task $L_{i}$.
@@
 This problem formulation constructs the learning process, such that the model learns how to adapt its parameters based on a few examples.
-This technique is popular in meta-learning [161], and is the foundation for many techniques (e.g, [162, 163, 164, 165, 166]).
+This technique is popular in meta-learning (Hospedales et al., 2021), and is the foundation for many techniques (e.g., Antoniou et al., 2018; Bertinetto et al., 2018; Finn et al., 2018; Nichol et al., 2018; Yoon et al., 2018).
 An alternative approach learns the parameters of an optimizer.
-An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples [167].
-The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer [168].
+An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples (Ravi and Larochelle, 2016).
+The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer (Li et al., 2017).
 The majority of domain generalization techniques exhibit many-to-many portability by construction.
-Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures [169].
-Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation [170], and others only require a few examples [171].
+Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures (Zhou et al., 2020).
+Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation (Xu et al., 2021), and others only require a few examples (Chen et al., 2022b).
 MAML and other meta-learning techniques exhibit high adaptability and are general enough to apply to shallow neural networks; thus, exhibiting a low model complexity.
-Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times [162].
-IV-F Few-shot Learning
-Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain [62].
-Definition IV.6
-(Few-shot Learning [62]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$.
-Few-shot learning is often characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1.
+Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times (Antoniou et al., 2018).
+Figure 2: A human-robot team conducting two tasks, where each task is characterized by a unique distribution across the seven workload components.
+These tasks were implemented as a part of an experimental design for developing a multi-dimensional workload estimation method for unknown tasks (Bhagat Smith et al., 2024).
+4.6 Few-shot Learning
+Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain (Wang et al., 2020b).
+Definition:.
+(Few-shot Learning) Given a set of source domains $DS$ = $D_{s_{1}},\dots,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},\dots,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$ (Wang et al., 2020b).
+Few-shot learning is characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1.
+This problem formulation is similar to domain generalization, but with additional complexity.
+A domain shift results from the novel physiological response that an unknown task can induce, but a label shift is a new workload value.
+An illustrative example demonstrating the primary difference is provided in Figure 2.
+The left task has high auditory and fine motor workload, but a low gross motor workload.
+The right task has a high gross motor workload, but a low auditory workload.
+Training a model to estimate gross motor workload on the left task will result in a model that likely performs poorly when estimating workload for the right task, as the range of gross motor workload values is too different.
+Few-shot learning techniques overcome the additional complexity of handling multiple distribution shifts simultaneously by requiring labeled target domain data.
 Prior work primarily focused on applying meta-learning techniques to solve few-shot learning problems.
 These techniques overlap significantly with meta-learning for domain generalization.
-Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML [141], Prototypical network [172], Conditional Neural Processes [173])).
-Meta-learning techniques broadly fall into three categories: 1) parameter-based 2) metric-based 3) probabilitstic.
-Parameter-based meta-learning algorithms were discussed in Section III.G.
-An illustrative example demonstrating the primary difference between meta-learning for domain generalization and few-shot learning is provided in Figure 2.
-Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared [64, 174, 175].
+Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML (Finn et al., 2017), Prototypical network (Snell et al., 2017), Conditional Neural Processes (Garnelo et al., 2018a))).
+Meta-learning techniques broadly fall into three categories: 1) parameter-based, 2) metric-based, and 3) probabilistic.
+Parameter-based meta-learning algorithms were discussed in the Domain Generalization section.
+Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared (Vinyals et al., 2016; Zhou et al., 2023a; Li et al., 2023).
 Two neural networks are trained to encode data, from the source and target domain, into a latent feature space.
@@
 Typically, multiple support set data points are passed through the networks simultaneously to enable the calculation of a range of complex distance metrics.
-(a) Conditional Neural Process
-(a) Conditional Neural Process
-(b) Latent Neural Process
-Figure 3: A comparison of two basic neural process types: (a) conditional neural processes and (b) latent neural process.
-Green boxes represent input and output values, blue boxes represent a neural network, yellow boxes represent aggregated latent features, and the pink circle represent an aggregation function.
-This figure is adapted from Figure 3 in [176].
-Matching networks perform a pairwise comparison between latent features using cosine similarity [64].
-Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space [172].
-Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks [172].
-Relation networks decompose the problem into separate modules: an embedding module and a relation module [177].
+Matching networks perform a pairwise comparison between latent features using cosine similarity (Vinyals et al., 2016).
+Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space (Snell et al., 2017).
+Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks (Snell et al., 2017).
+Relation networks decompose the problem into separate modules: an embedding module and a relation module (Sung et al., 2018).
 The embedding module is a neural network designed for feature extraction and the relation module is a non-linear classifier trained on the latent feature representation.
-Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class [178].
+Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class (Simon et al., 2020).
 This technique maximizes the distance between subspaces by using a Grassmannian geometry-based objective function.
-Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g, [179, 180, 181]).
-Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process [182])) and a neural network.
-Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets [176].
-Probabilistic meta-learning techniques fall into two broad categories: Neural Processes [183], and Deep Kernel Learning [184].
-Neural processes pass the data through a neural network, parameterized by $\theta$, to learn a latent representation.
-Each support set data point is passed through an aggregation function, $a$, to calculate the latent feature, $z$.
-This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes [176]).
-Another neural network, parameterized by $\phi$, uses $z$ and one or more data points (e.g., one-shot vs.
+Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g., Kim et al., 2019; Schonfeld et al., 2019; Zhang et al., 2019).
+Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process (Rasmussen, 2003)) and a neural network.
+Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets (Jha et al., 2022).
+Probabilistic meta-learning techniques fall into two broad categories: Neural Processes (Garnelo et al., 2018b), and Deep Kernel Learning (Wilson et al., 2016).
+Neural processes pass the data through a neural network to learn a latent representation.
+Each support set data point is passed through an aggregation function to calculate the latent feature.
+This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes (Jha et al., 2022)).
+Another neural network uses that latent representation and one or more data points (e.g., one-shot vs.
 few-shot) from the target domain to learn the parameters of a Gaussian Process.
-There are two main neural process types: latent neural processes, and conditional neural processes, as shown in Figure 3.
-Latent neural processes construct a stochastic variable using latent features, which typically takes the form of a Gaussian distribution parameterized by $\mu$ and $\sigma$ [183].
-However, by modeling these correlations the underlying predictive model becomes intractable and must be approximated (e.g., Monte Carlo Sampling [185], Variational Inference [183]).
-Conditional neural processes do not construct a stochastic variable using latent features [173], instead they make the assumption the stochastic process can be fully described by these encoded values (i.e., factorized Gaussian distribution [176]).
-This assumption makes the underlying predictive model analytically tractable, but ignores any correlations that may exist between latent features.
-Several neural network architectural components have been incorporated into conditional neural processes with varying degrees of success (e.g., [186, 187, 188, 189, 190, 191, 192, 193]).
-Deep kernel learning techniques takes an alternative by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, effectively training the network to learn a custom kernel function for that Gaussian process [184].
-Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks [194].
-This technique learns over a large number of small, but related, examples that train the underlying machine learning model to approximate a prior distribution to transfer knowledge between learning tasks.
+There are two main neural process types: latent neural processes (Garnelo et al., 2018b) and conditional neural processes (Garnelo et al., 2018a).
+Conditional neural processes have received the most attention, where many techniques incorporate different neural network architectural components with varying degrees of success (e.g., Gu et al., 2023a; Holderrieth et al., 2021; Bruinsma et al., 2023; Kim et al., 2022; Wang et al., 2022; Ye and Yao, 2022; Dexheimer and Davison, 2023; Pandey and Yu, 2023)).
+Deep kernel learning techniques take an alternative approach by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, training the network to learn a custom kernel function for that Gaussian process (Wilson et al., 2016).
+Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks (Patacchiola et al., 2020).
+This technique learns over a large number of small, but related, examples that train the machine learning model to approximate a prior distribution to transfer knowledge between learning tasks.
 Deep kernel transfer quantifies uncertainty by maintaining a distribution over parameters of the neural network, which can become computationally expensive for larger networks.
-Pólya-Gamma augmentation and softmax approximation were explored to overcome this limitation by maintaining a distribution over functions [195].
-Deep mean function learning has also been applied as a meta-learning technique, where a neural network is used to learn a custom mean function for a Gaussian process [196].
-These and similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., [156, 197, 198, 199, 200]).
-However, recent results demonstrate that the original deep kernel learning algorithm tends to produce unreliable uncertainty estimates in some practical application domains [201].
-Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (e.g., [201][202])
+Similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., Rothfuss et al., 2023; Wei et al., 2023; Bånkestad et al., 2023; Maraval et al., 2024; Iwata and Kumagai, 2024; Lee et al., 2023).
+However, recent results demonstrate that many deep kernel learning algorithms tend to produce unreliable uncertainty estimates in some application domains (van Amersfoort et al., 2021).
+Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (Liu et al., 2023).
 Few-shot learning techniques exhibit many-to-many portability and high adaptability.
-Most few-shot learning techniques were developed using computer vision benchmarks (e.g., [203, 172, 64]), which requires deep learning architectures.
-However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model [11].
+Most few-shot learning techniques were developed using computer vision benchmarks (e.g., Garcia and Bruna, 2018; Snell et al., 2017; Vinyals et al., 2016), which requires deep learning architectures.
+However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model (Tian et al., 2020).
 Therefore, it stands to reason that model complexity is directly related to the learning task’s complexity.
@@
 Therefore, meta-learning presents a promising option for develop workload estimation for unknown tasks solutions and merit future investigation.
-V Discussion
-Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are of particular importance to each task.
-Understanding the distribution across workload components is critical to adapting interactions and task allocations to maximize the performance of a human-robot team.
+5 Discussion
+Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are particularly important for each task.
+Understanding the distribution across workload components is critical to adapting robot behavior to maximize the human-robot team’s performance.
 Quantifying uncertainty helps enumerate this distribution by contextualizing component-specific workload estimates, as components may rely on overlapping metrics (e.g., fine-motor and tactile).
-Domain adaptation and transfer learning are popular methods, but require high model complexity to achieve desired performance levels.
-These methods are also best suited to one-to-one portability problems with low adaptability.
-Specifically, these techniques are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains and where all data is gathered prior to training.
-These techniques do not adapt machine learning models in the moment; thus, are not viable for performing workload estimation for unknown tasks in real world application domains.
-Test-time adaptation overcomes this particular constraint, but still exhibits one-to-one portability and low adaptability.
+Viewing candidate machine learning methods through the three developed criteria (i.e., portability, model complexity, and adaptability) facilitates discussing for how each candidate addresses the core concerns of estimating human workload for unknown tasks (see Evaluation Criteria section).
+Domain adaptation and transfer learning are popular methods, but are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains, and where all data is gathered prior to training (i.e., one-to-one portability and low adaptability).
+These methods are not well suited for estimating a human’s workload for unknown tasks, as it is impractical to gather a significant amount of data for new tasks (i.e., target domain) prior to training the model.
+Additionally, these techniques do not adapt machine learning models in real time.
+Test-time adaptation overcomes this particular constraint, but exhibits one-to-one portability and low adaptability.
 Existing test-time adaptation techniques applied to smaller neural networks may exhibit higher adaptability, but further investigation is required.
+Regardless, none of these techniques are well suited for estimating human workload in the presence of unknown tasks.
 Continual learning has been frequently proposed as a potential solution to non-IID machine learning problems within HRI, but are not applicable to workload estimation for unknown tasks due to their poor portability.
@@
 Consider an application domain with $N$ tasks.
-Continual learning must learn $N!$ task-pair independent updates, as the tasks’ order influences the learning process.
+Continual learning must learn $N!$ task-pair independent updates, as the tasks’ ordering influences the learning process.
 Domain generalization and few-shot learning techniques only learn $N$ updates, as they can trace each task to a single common prior.
 These additional portability-related constraints make continual learning a subpar option.
-Domain generalization and few-shot learning exhibit desirable properties for all three criteria: high adaptability, one-to-one portability, and low model complexity.
-These techniques are explicitly designed to update a machine learning model based upon limited and highly variable data, and are flexible enough change the model complexity based upon the learning task’s complexity.
-Multiple domain shifts from a common prior is the most directly related metaphor for the variability seen in human behavior; thus, domain generalization is the most suited for scenarios where there is no label shift (e.g., workload estimation).
-Few-shot learning may be more applicable to scenarios where labels change between learning tasks or application domains (i.e., task recognition, affect recognition).
-The presented meta-learning techniques are general enough to apply to either machine learning problem.
-Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem [1, 204, 21].
-Only a limited number of techniques (e.g., probablistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks.
-Probablistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
-A key research question is if existing workload estimation datasets capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques can accurately estimate workload for unknown tasks.
-It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial.
-These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID and non-IID tasks that these techniques can handle.
-Understanding how probablistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real world viability.
-Additionally, analyzing how the differences between tasks, human-robot interactions, robot capabilities further informs how these techniques can be successfully deployed.
-Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown tasks; thus, merit further investigation.
-VI Conclusion
-Developing models that can estimate the human’s internal state is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments.
+Domain generalization and few-shot learning exhibit desirable properties for all criteria: high adaptability, one-to-one portability, and low model complexity.
+These techniques are designed to update a machine learning model based on limited and highly variable data, and are flexible enough to change the model complexity based on the learning task’s complexity.
+Multiple domain shifts from a common prior is the most directly related metaphor for task-based variability, as the contribution of each physiological signal to each workload component and the contribution of each workload component will vary between tasks.
+Therefore, domain generalization is the most suited for workload estimation for unknown tasks when there is no label shift.
+Few-shot learning is more applicable to scenarios where there are significant changes to the range of workload values between tasks (i.e., label shift).
+Meta-learning techniques sit at the intersection of these two categories and are general enough to apply to either learning task.
+Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem (Heard et al., 2019; Bhagat Smith et al., 2023, 2022).
+Only a limited number of techniques (e.g., probabilistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks.
+Probabilistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
+Existing workload estimation datasets frequently struggle to capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques’ ability to accurately estimate workload for novel tasks can be assessed.
+It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial (Bhagat Smith et al., 2024).
+These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID (i.e.
+known) and non-IID tasks (i.e., unknown) that these techniques can handle.
+Understanding how probabilistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real-world viability.
+Additionally, analyzing how the differences between tasks, human-robot interactions, and robot capabilities further informs how these techniques can be successfully deployed.
+Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown novel tasks; thus, merit further investigation.
+6 Conclusion
+Developing models that can estimate the human’s workload is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments.
 The variability of these environments is characterized by varying human-robot teaming dynamics, diverse tasks executions, and different operating conditions for the team itself.
-A robot seeking to adapt its behavior to best assist its human teammate must accurate the human’s internal state accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible.
-Non-IID machine learning, particular domain generalization and few-shot learning, hold promise in developing models that can account for these varying dynamics.
-Acknowledgments
-The graduate students have been supported by National Aeronautics and Space Administration through University Space Research Association aware 904186092.
-The contents are those of the authors and do not represent the official views of, nor an endorsement, by the National Aeronautics and Space Administration.
-The presented work was partially supported by ONR grants N00024-20-F-8705 and N00014-21-1-2052.
-The views, opinions, and findings expressed are those of the authors and are not to be interpreted as representing the official views or policies of the Department of Defense or the U.S.
-Government.
+A robot seeking to adapt its behavior to best assist its human teammate must estimate the human’s workload accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible.
+This manuscript developed three criteria (i.e., portability, model complexity, and adaptability) and presented an analysis of relevant non-IID machine learning techniques.
+Numerous non-IID machine learning techniques exist, so this manuscript assessed how each technique’s machine learning considerations interact with the unique characteristics of HRI datasets (e.g., workload estimation for unknown tasks).
+A deeper understanding of these interactions highlight the specific requirements for a particular application, which is useful for guiding future empirical research.
+This manuscript’s analysis argues that domain generalization and few-shot learning techniques hold promise in developing models that can estimate a human’s workload for unknown tasks.